# Backdoor Attack in Prompt-Based Continual Learning

### v1

@@ Abstract
Prompt-based approaches offer a cutting-edge solution to data privacy issues in continual learning, particularly in scenarios involving multiple data suppliers where long-term storage of private user data is prohibited. Despite delivering state-of-the-art performance, its impressive remembering capability can become a double-edged sword, raising security concerns as it might inadvertently retain poisoned knowledge injected during learning from private user data. Following this insight, in this paper, we expose continual learning to a potential threat: backdoor attack, which drives the model to follow a desired adversarial target whenever a specific trigger is present while still performing normally on clean samples. We highlight three critical challenges in executing backdoor attacks on incremental learners and propose corresponding solutions: (1) Transferability: We employ a surrogate dataset and manipulate prompt selection to transfer backdoor knowledge to data from other suppliers; (2) Resiliency: We simulate static and dynamic states of the victim to ensure the backdoor trigger remains robust during intense incremental learning processes; and (3) Authenticity: We apply binary cross-entropy loss as an anti-cheating factor to prevent the backdoor trigger from devolving into adversarial noise. Extensive experiments across various benchmark datasets and continual learners validate our continual backdoor framework, achieving up to $100\%$ attack success rate, with further ablation studies confirming our contributions’ effectiveness.
@@ 1 Introduction
Nonetheless, such promising results can inadvertently become vulnerabilities, exposing CL to security threats. Indeed, while CL methods effectively address catastrophic forgetting by preserving and incorporating previously acquired knowledge, they may also unwittingly retain knowledge compromised by adversarial actions. These threats become even more formidable in the multi-data supplier scenario of prompt-based approaches, where the supplied data might contain hidden harmful information.
@@
Challenges Despite holding such potential danger for CL, extending backdoor attacks to the incremental setting is non-trivial. Firstly, in the multi-supplier setting where the victim gathers data from different sources, the attacker lacks information about the actual data distribution used to train the victim model. Consequently, generalizing backdoor knowledge to be transferable to unknown data poses the first challenge that our continual backdoor approach must confront. The second challenge arises from the vulnerability of backdoor attacks during fine-tuning. Recent studies [44, 35] have highlighted the tendency for backdoor knowledge to be removed when the victim fine-tunes the poisoned model on a small and clean dataset. This issue is exacerbated in continual learning, where the victim model undergoes incremental training as new data from various sources arrive. The final challenge involves the backdoor trigger’s proneness to turn into adversarial noise. Huynh et al. [18] observed that the trigger, when optimized using a surrogate model, may transform into an adversarial perturbation, driving the clean model to follow desired adversarial targets even in the absence of any prior backdoor attacks. Since conventional adversarial defenses can mitigate such adversarial noise, preempting this behavior is crucial to strengthen the resilience of the backdoor trigger.
Contributions In response to these shortcomings, we propose a continual backdoor framework that satisfies three key properties: transferability to unknown data, resilience to incremental learning procedures, and authenticity to avoid becoming adversarial noise. Initially, we leverage the natural label mapping characteristic of visual prompting, thereby approaching the data poisoning issue from the perspective of prompt selection. This approach allows our backdoor trigger to be generalized to any victim data distribution. Next, we robustify the backdoor trigger by aligning the optimization process with the continuously changing states of the incremental learner, thus ensuring the effectiveness of the backdoor trigger when the model is trained on new incoming clean data. Finally, we reconsider the choice of loss function for trigger optimization. We observe that the commonly used softmax function with cross-entropy introduces bias towards the target class, pushing its score excessively high and leading to the adversarial noise problem. Building on this observation, we propose adopting binary cross-entropy (BCE) with sigmoid function to mitigate this issue, thereby eliminating the dependency of trigger optimization on other classes and preventing cheating behavior.
@@
Organization The paper is organized as follows. Section 2 provides a brief overview of continual learning and prompt-based continual learning. In Section 3, we introduce the continual backdoor threat model, discuss backdoor challenges, and propose our prompt-based continual backdoor AOP framework. Section 4 empirically verifies the effectiveness of our AOP framework against various prompt-based incremental learners. Finally, Section 5 concludes the paper. Additional related work, discussions, and experiments are included in the supplementary material.
@@ 2 Background
Continual learning In continual learning scenarios, the model undergoes a sequential presentation of tasks ${\mathcal{D}_{1},...\mathcal{D}_{T}}$. Each task corresponds to distinct subsets of tuples $\mathcal{D}_{t}=\{\boldsymbol{x}_{t}^{i},\boldsymbol{y}_{t}^{i}\}_{i=1}^{i=n_{t}}$, where $\boldsymbol{x}_{t}^{i}\in\mathcal{X}^{t}$ is the input sample, $\boldsymbol{y}_{t}^{i}\in\mathcal{Y}^{t}$ is the corresponding label, and $n_{t}$ is the number of samples for task $t$. It is important to note that each class is exclusively associated with a single task [7, 3], meaning that $\mathcal{Y}^{t}$ and $\mathcal{Y}^{t^{\prime}}$ are disjoint, and data from prior tasks become inaccessible during the training of subsequent tasks [45, 40]. The objective of continual learning is to continuously acquire the capability to classify newly introduced classes while maintaining proficiency on previously learned ones in a single model $f:\mathcal{X}\rightarrow\mathcal{Y}$. In this paper, and in prompt-based methods [45, 57, 58, 55, 40], $f$ represents the pre-trained Vision Transformer (ViT) encoder. Additionally, $\phi$ is employed as the shared classification head, and $\phi_{t}$ is the classifier corresponding to classes specific to the given task $t$.
Prompt-based continual learning We provide a concise overview of L2P [58], which stands as the first work that integrates prompts into the context of continual learning. L2P introduces a prompt pool comprising learnable prompts and their corresponding keys $\left\{\left(\boldsymbol{k}_{1},\boldsymbol{p}_{1}\right),\left(\boldsymbol{k}_{2},\boldsymbol{p}_{2}\right),\cdots,\left(\boldsymbol{k}_{n_{p}},\boldsymbol{p}_{n_{p}}\right)\right\}$ where $n_{p}$ is total number of prompts. These prompts are then combined with image features and fed into a pre-trained ViT, instructing the model to perform classification. Prompts are queried in an instance-wise manner using the top-$K$ cosine similarity $\gamma\left(q(\boldsymbol{x}),\boldsymbol{k}_{i}\right)$ between the keys and the query function $q(\boldsymbol{x})=f(\boldsymbol{x})[0,:]$. Subsequent prompt-based methods are designed based on L2P, each featuring prompt utility and optimization modifications. A brief explanation of these methods is in Appendix A.
@@ 3 Backdoor Attack on Prompt-based Continual Learning (AOP)
We first outline the threat model and introduce key notations in Section 3.1. Then, we highlight the challenges when executing a backdoor attack against prompt-based incremental learners in Section 3.2. Building upon these considerations, we delineate the three primary components of AOP across Sections 3.3-3.5. A comprehensive overview and the end-to-end algorithm is in Appendix B.
@@ 3.1 Threat Model and Notations
Continual learning protocols We consider the class-incremental learning (CIL) setting in prompt-based continual learning [57, 58]. In CIL, training data for incremental tasks $\mathcal{D}_{t}$ arrive incrementally in a discrete manner. Each task consists of data for new $M$ classes that have not been learned by the model before. Formally, each task $\mathcal{D}_{t}={\{\mathcal{D}_{m,t}\}_{m=1}^{M}}$ with each class $\mathcal{D}_{m,t}={\{\boldsymbol{x}_{i}^{m,t},y_{i}^{m,t}\}}^{n_{m,t}}_{i=1}$ comprises input samples $\boldsymbol{x}_{i}^{m,t}\in\mathcal{X}$ and their corresponding labels $y_{i}^{m,t}=c_{m,t}\in\mathcal{Y}$, where $n_{m,t}$ represents the number of training samples for the corresponding class. In CIL, the learner is required to perform classification across all classes encountered up to task $T$ without being provided with explicit task labels during inference. Data for different classes $m$ and $m^{\prime}$ are gathered from different suppliers. To ease the ensuing presentation, the index $t$ is omitted unless noted otherwise.
Backdoor attack protocols Let the attacker be the data supplier for class $m$ with labels $c_{m}$. The attacker’s goal is to poison the supplying dataset with a small amount of trigger-injected samples, such that any data from any classes if manipulated with the backdoor trigger, will be misclassified as $c_{m}$ by the resulting incremental victim model when performing inference at any time $t$. An example of a triggered image is given in Figure 4.
Consider $\mathcal{D}_{m}=\left\{\left(\boldsymbol{x}_{i},y_{i}\right)\right\}_{i=1}^{n_{m}}$ as the benign training set of class $m$. The adversary then learns to generate the poisoned dataset $\mathcal{D}_{p}$. Specifically, $\mathcal{D}_{p}$ consists of two parts: a modified version of a selected subset (denoted as $\mathcal{D}_{s}$) of $\mathcal{D}_{m}$ and the remaining benign samples. Thus, $\mathcal{D}_{p}=\mathcal{D}_{b}\cup\mathcal{D}_{c}$, where $c_{m}$ is the adversary target label, $\mathcal{D}_{c}=\mathcal{D}_{m}\backslash\mathcal{D}_{s}$, $\mathcal{D}_{b}=\left\{\left(\boldsymbol{x}^{\prime},c_{m}\right)\mid\boldsymbol{x}^{\prime}=G(\boldsymbol{x}),(\boldsymbol{x},c_{m})\in\mathcal{D}_{s}\right\}$, $\gamma\triangleq\frac{\left|\mathcal{D}_{s}\right|}{|\mathcal{D}_{m}|}$ is the poisoning rate, and $G:\mathcal{X}\rightarrow\mathcal{X}$ is an adversary-specified poisoned image generator. We follow [46, 29] and formulate $G(\boldsymbol{x})=\boldsymbol{x}+\boldsymbol{\delta}$, where the perturbation $\boldsymbol{\delta}$ has a bounded $\ell_{p}$-norm.
@@
3.2 Three Challenges When Backdooring CL
We outline three challenges encountered when executing backdoor attacks against continual learners. To generate a poisoned dataset, the adversary optimizes the backdoor trigger, necessitating the appearance of the training data, learner, and training criterion. However, in accordance with our threat model, none of these are accessible.
The first challenge, as outlined in the introduction, arises from the lack of knowledge about the victim’s training data. Given that control is limited to the supplied data, which also represents the target class, prior research [64] suggests utilizing a public dataset (e.g., Tiny-ImageNet) as a surrogate training dataset. In this study, we explore the utilization of surrogate datasets in the context of prompt tuning.
Secondly, the adversary lacks information about the training learner and procedure, making it difficult to design backdoor knowledge that can withstand the incremental learning process. Despite impressive memory capabilities, continual learning methods have not yet fully matched the performance levels of joint training, and recent works [40] are still exploring ways to further avoid catastrophic forgetting. This issue also affects backdoor attacks, leading to a degradation in attack performance over time. Therefore, creating a surrogate learner that helps the trigger endure the incremental learning process is our second challenge.
Lastly, a backdoor attack entails poisoning the training dataset to induce the model to malfunction when presented with specific trigger samples while maintaining normal performance on clean data. Huynh et al. [18] observe that this objective can be achieved even without any poisoning during training. This trigger, akin to adversarial noise, can deceive the classifier during inference, irrespective of whether data poisoning occurred during training, thereby counteracting the primary objective of the backdoor attack. Moreover, such adversarial noise can be mitigated by employing standard adversarial defenses. Consequently, preventing the generation of adversarial noise poses an additional challenge when optimizing the trigger.
@@ 3.3 Prompt Selection, Label Mapping, and Transferability
The core of prompt-based continual learning methods lies in the prompt pool and the prompt selection strategy. Specifically, the most relevant prompts are queried in an instance-wise manner and then concatenated with the sample to optimally guide the model in performing classification. We leverage this fundamental mechanism of the prompt-based approach to reframe the backdooring problem as one of manipulating prompt selections. As in Figures 1(a) and 1(b), we aim to ensure that triggered samples are directed to select specific backdoor prompts, thereby causing the model to misclassify these backdoor-prompted samples into the desired class.
@@ 3.4 Static-dynamic Trigger Optimization
Since we lack information about the victim’s continual model, we use $\mathcal{D}_{\text{surrogate}}$ to train a surrogate incremental learner and simulate the continual learning pipeline. We then optimize the backdoor trigger $\boldsymbol{\delta}$ based on this surrogate incremental model. Specifically, we employ the surrogate learner with two states: a static state that reflects how prompts learn label mappings between the source and target datasets, and a dynamic state that reflects the continuous learning procedure of the victim model. Formally, our static-dynamic trigger optimization involves the following four stages:
Figure 2: The AOP procedure begins by selecting a surrogate dataset, which is then divided into two subsets: $\mathcal{D}_{\text{static}}$ and $\mathcal{D}_{\text{dynamic}}$. In stage (1), $\mathcal{D}_{\text{static}}$ is employed to establish a static surrogate learner along with prompts. Following this, in stage (2), the trigger optimization process takes place based on this initial model. Next, in stage (3), the learner is updated from $\mathcal{D}_{\text{dynamic}}$, which serves as a transition between stages. Finally, in stage (4), the trigger is fine-tuned, with the prompt being updated periodically throughout the optimization process.
@@
$\begin{array}[]{r}\min_{\mathbf{P},\mathbf{K},\phi}\mathcal{L}\left(\phi\left(f\left(\boldsymbol{x};\boldsymbol{P}\right)\right),y\right)+\lambda\sum_{\mathbf{K}_{\boldsymbol{x}}}\gamma\left(q(\boldsymbol{x}),\boldsymbol{k}_{i}\right).\\ \end{array}$ (1)
@@
(4) Dynamic surrogate stage In this stage, we aim to acquaint the backdoor trigger with the continuously updated prompts resulting from the continual learning process. This dynamic stage entails fine-tuning the prompt components for one epoch, as in Stage (3), following several iterations of optimization of the trigger with equation (2). This iterative process is repeated for multiple rounds to enhance the resilience of the backdoor trigger against the continual learning process.
After optimizing the trigger through the aforementioned four stages, the optimized trigger $\boldsymbol{\delta}^{*}$ is used to poison a small portion of $\mathcal{D}_{m}$, which is then released to the victim learner. Summarization of AOP is in Figure 2 and Appendix B.

### v2

# Attack On Prompt: Backdoor Attack in Prompt-Based Continual Learning

@@ Abstract
Prompt-based approaches offer a cutting-edge solution to data privacy issues in continual learning, particularly in scenarios involving multiple data suppliers where long-term storage of private user data is prohibited. Despite delivering state-of-the-art performance, its impressive remembering capability can become a double-edged sword, raising security concerns as it might inadvertently retain poisoned knowledge injected during learning from private user data. Following this insight, in this paper, we expose continual learning to a potential threat: backdoor attack, which drives the model to follow a desired adversarial target whenever a specific trigger is present while still performing normally on clean samples. We highlight three critical challenges in executing backdoor attacks on incremental learners and propose corresponding solutions: (1) Transferability: We employ a surrogate dataset and manipulate prompt selection to transfer backdoor knowledge to data from other suppliers; (2) Resiliency: We simulate static and dynamic states of the victim to ensure the backdoor trigger remains robust during intense incremental learning processes; and (3) Authenticity: We apply binary cross-entropy loss as an anti-cheating factor to prevent the backdoor trigger from devolving into adversarial noise. Extensive experiments across various benchmark datasets and continual learners validate our continual backdoor framework, with further ablation studies confirming our contributions’ effectiveness.
@@ 1 Introduction
Nonetheless, such promising results can inadvertently become vulnerabilities, exposing CL to security threats. Indeed, while CL methods effectively address catastrophic forgetting by preserving and incorporating previously acquired knowledge, they may also unwittingly retain knowledge compromised by adversarial actions. These threats become even more formidable in the multi-data supplier scenario of prompt-based approaches, where the supplied data might contain hidden harmful information, as shown in Figure 1.
Figure 1: Multi data supplier scenario in prompt-based continual learning, with one supplier acting as an adversarial attacker.
@@
Challenges. Despite holding such potential danger for CL, extending backdoor attacks to the incremental setting is non-trivial. Firstly, in the multi-supplier setting where the victim gathers data from different sources, the attacker lacks information about the actual data distribution used to train the victim model. Consequently, generalizing backdoor knowledge to be transferable to unknown data poses the first challenge that our continual backdoor approach must confront. The second challenge arises from the vulnerability of backdoor attacks during fine-tuning. Recent studies [44, 35] have highlighted the tendency for backdoor knowledge to be removed when the victim fine-tunes the poisoned model on a small and clean dataset. This issue is exacerbated in continual learning, where the victim model undergoes incremental training as new data from various sources arrive. The final challenge involves the backdoor trigger’s proneness to turn into adversarial noise. Huynh et al. [18] observed that the trigger, when optimized using a surrogate model, may transform into an adversarial perturbation, driving the clean model to follow desired adversarial targets even in the absence of any prior backdoor attacks. Since conventional adversarial defenses can mitigate such adversarial noise, preempting this behavior is crucial to strengthen the resilience of the backdoor trigger.
Contributions. In response to these shortcomings, we propose a continual backdoor framework that satisfies three key properties: transferability to unknown data, resilience to incremental learning procedures, and authenticity to avoid becoming adversarial noise. Initially, we leverage the natural label mapping characteristic of visual prompting, thereby approaching the data poisoning issue from the perspective of prompt selection. This approach allows our backdoor trigger to be generalized to any victim data distribution. Next, we robustify the backdoor trigger by aligning the optimization process with the continuously changing states of the incremental learner, thus ensuring the effectiveness of the backdoor trigger when the model is trained on new incoming clean data. Finally, we reconsider the choice of loss function for trigger optimization. We observe that the commonly used softmax function with cross-entropy introduces bias towards the target class, pushing its score excessively high and leading to the adversarial noise problem. Building on this observation, we propose adopting binary cross-entropy (BCE) with sigmoid function to mitigate this issue, thereby eliminating the dependency of trigger optimization on other classes and preventing cheating behavior.
@@ 2 Background
Continual learning. In continual learning scenarios, the model undergoes a sequential presentation of tasks ${\mathcal{D}_{1},...\mathcal{D}_{T}}$. Each task corresponds to distinct subsets of tuples $\mathcal{D}_{t}=\{\boldsymbol{x}_{t}^{i},\boldsymbol{y}_{t}^{i}\}_{i=1}^{i=n_{t}}$, where $\boldsymbol{x}_{t}^{i}\in\mathcal{X}^{t}$ is the input sample, $\boldsymbol{y}_{t}^{i}\in\mathcal{Y}^{t}$ is the corresponding label, and $n_{t}$ is the number of samples for task $t$. It is important to note that data from prior tasks become inaccessible during the training of subsequent tasks [45, 40]. The objective of continual learning is to continuously acquire the capability to classify newly introduced classes while maintaining proficiency on previously learned ones in a single model $f:\mathcal{X}\rightarrow\mathcal{Y}$. In this paper, and in prompt-based methods [45, 57, 58, 55, 40], $f$ represents the pre-trained Vision Transformer (ViT) encoder. Additionally, $\phi$ is employed as the shared classification head, and $\phi_{t}$ is the classifier corresponding to classes specific to the given task $t$.
Prompt-based continual learning. We provide a concise overview of L2P [58], which stands as the first work that integrates prompts into the context of continual learning. L2P introduces a prompt pool comprising learnable prompts and their corresponding keys $\left\{\left(\boldsymbol{k}_{1},\boldsymbol{p}_{1}\right),\left(\boldsymbol{k}_{2},\boldsymbol{p}_{2}\right),\cdots,\left(\boldsymbol{k}_{n_{p}},\boldsymbol{p}_{n_{p}}\right)\right\}$ where $n_{p}$ is total number of prompts. These prompts are then combined with image features and fed into the pre-trained ViT $f$, instructing the model to perform classification. Prompts are queried in an instance-wise manner using the top-$K$ cosine similarity $\gamma\left(q(\boldsymbol{x}),\boldsymbol{k}_{i}\right)$ between the keys and the query function $q(\boldsymbol{x})=f(\boldsymbol{x})[0,:]$. Subsequent prompt-based methods are designed based on L2P, each featuring prompt utility and optimization modifications. A brief explanation of these methods can be found in the supplementary material.
@@ 3 Backdoor Attack on Prompt-based Continual Learning (AOP)
We first outline the threat model and introduce key notations in Section 3.1. We then delineate the three primary components of AOP across Sections 3.2-3.4. A comprehensive overview and the end-to-end algorithm is in the supplementary material.
@@ 3.1 Threat Model and Notations
Continual learning protocols. We consider the class-incremental learning (CIL) setting in prompt-based continual learning [57, 58]. In CIL, training data for incremental tasks $\mathcal{D}_{t}$ arrive incrementally in a discrete manner. Each task consists of data for new $M$ classes that have not been learned by the model before. Formally, each task $\mathcal{D}_{t}={\{\mathcal{D}_{m,t}\}_{m=1}^{M}}$ with each class $\mathcal{D}_{m,t}={\{\boldsymbol{x}_{i}^{m,t},y_{i}^{m,t}\}}^{n_{m,t}}_{i=1}$ comprises input samples $\boldsymbol{x}_{i}^{m,t}\in\mathcal{X}$ and their corresponding labels $y_{i}^{m,t}=c_{m,t}\in\mathcal{Y}$, where $n_{m,t}$ represents the number of training samples for the corresponding class. In CIL, the learner is required to perform classification across all classes encountered up to task $T$ without being provided with explicit task labels during inference. Data for different classes $m$ and $m^{\prime}$ are gathered from different suppliers. To ease the ensuing presentation, the index $t$ is omitted unless noted otherwise.
Backdoor attack protocols. Let the attacker be the data supplier for class $m$ with labels $c_{m}$. The attacker’s goal is to poison the supplying dataset with a small amount of trigger-injected samples, such that any data from any classes if manipulated with the backdoor trigger, will be misclassified as $c_{m}$ by the resulting incremental victim model when performing inference at any time $t$. An example of a triggered image is given in the supplementary material.
Consider $\mathcal{D}_{m}=\left\{\left(\boldsymbol{x}_{i},y_{i}\right)\right\}_{i=1}^{n_{m}}$ as the benign training set of class $m$. The adversary then learns to generate the poisoned dataset $\mathcal{D}_{p}$. Specifically, $\mathcal{D}_{p}$ consists of two parts: a modified version of a selected subset (denoted as $\mathcal{D}_{s}$) of $\mathcal{D}_{m}$ and the remaining benign samples. Thus, $\mathcal{D}_{p}=\mathcal{D}_{c}\cup\mathcal{D}_{b}$, $\mathcal{D}_{c}=\mathcal{D}_{m}\backslash\mathcal{D}_{s}$, $\mathcal{D}_{b}=\left\{\left(\boldsymbol{x}^{\prime},c_{m}\right)\mid\boldsymbol{x}^{\prime}=G(\boldsymbol{x}),(\boldsymbol{x},c_{m})\in\mathcal{D}_{s}\right\}$, where $c_{m}$ is the adversary target label, $\gamma\triangleq\frac{\left|\mathcal{D}_{s}\right|}{|\mathcal{D}_{m}|}$ is the poisoning rate, and $G:\mathcal{X}\rightarrow\mathcal{X}$ is an adversary-specified poisoned image generator. We follow [46, 29] and formulate $G(\boldsymbol{x})=\boldsymbol{x}+\boldsymbol{\delta}$, where the perturbation $\boldsymbol{\delta}$ has a bounded $\ell_{p}$-norm.
@@ 3.2 Prompt Selection, Label Mapping, and Transferability
The core of prompt-based continual learning methods lies in the prompt pool and the prompt selection strategy. Specifically, the most relevant prompts are queried in an instance-wise manner and then concatenated with the sample to optimally guide the model in performing classification. We leverage this fundamental mechanism of the prompt-based approach to reframe the backdooring problem as one of manipulating prompt selections. As in Figures 2(a) and 2(b), we aim to ensure that triggered samples are directed to select specific backdoor prompts, thereby causing the model to misclassify these backdoor-prompted samples into the desired class.
@@ 3.3 Static-dynamic Trigger Optimization
Since we lack information about the victim’s continual model, we use $\mathcal{D}_{\text{surrogate}}$ to train a surrogate incremental learner. We then optimize the backdoor trigger $\boldsymbol{\delta}$ based on this surrogate incremental model. Specifically, we employ the surrogate learner with two states: a static state that reflects how prompts learn label mappings between the source and target datasets, and a dynamic state that reflects the continuous learning procedure of the victim model. Formally, our static-dynamic trigger optimization involves the following four stages:
@@
$\min_{\mathbf{P},\mathbf{K},\phi}\mathcal{L}\left(\phi\left(f\left(\boldsymbol{x};\boldsymbol{P}\right)\right),y\right)+\lambda\sum_{\mathbf{K}_{\boldsymbol{x}}}\gamma\left(q(\boldsymbol{x}),\boldsymbol{k}_{i}\right).$ (1)
@@
Figure 3: AOP framework. The backdoor trigger and prompt pool are updated using the static-dynamic strategy. Clean and poisoned data are mapped to corresponding prompts, guiding the pretrained model to behave normally on clean inputs while misclassifying triggered inputs according to the adversary’s target.
@@
(4) Dynamic surrogate stage In this stage, we aim to acquaint the backdoor trigger with the continuously updated prompts in continual learning. This dynamic stage entails fine-tuning the prompt components for one epoch, as in Stage (3), following several iterations of optimization of the trigger with equation (2). This iterative process is repeated for multiple rounds to enhance the resilience of the backdoor trigger against the continual learning process.
After optimizing the trigger through the aforementioned four stages, the optimized trigger $\boldsymbol{\delta}^{*}$ is used to poison a small portion of $\mathcal{D}_{m}$, which is then released to the victim learner. Summarization of AOP is in Figure 3 and further details are in the supplementary material.